# Holmes: Towards Effective and Harmless Model Ownership Verification to Personalized Large Vision Models via Decoupling Common Features

### v1

@@ III-A Preliminaries
where $\mathcal{L}(\cdot)$ is the loss function ($i.e.$, cross-entropy), and $\bm{\theta}$ represents the parameters of the foundation model $F$. This fine-tuning process adapts $F$ to the specific task defined by $\mathcal{D}$, rather than trains the model from scratch.
@@
For a model ownership verification method to be effective in practice, it must satisfy two key requirements: effectiveness and harmlessness. Effectiveness means that the defense must accurately determine whether the suspicious model has been stolen from the victim, regardless of the model stealing technique employed. At the same time, it should not misclassify independently trained or fine-tuned benign models as stolen models. Harmlessness ensures that the defense does not introduce additional security risks, such as backdoors. Furthermore, if the method modifies the training set, harmlessness implies that models trained or fine-tuned with the modified dataset should exhibit behavior similar to those trained with the original dataset. A stricter standard of harmlessness requires that the defense does not modify the training or fine-tuning process of the victim model at all.
@@
Threat Model. In this paper, we consider the black-box setting for model ownership verification. Specifically, we assume that the defenders can only query and obtain the predicted probabilities (logit outputs) from the suspicious model, with no access to: model source files or parameters, intermediate computational results ($e.g.$, gradients), or any information about the potential model stealing process.
III-B Revisiting Harmless Model Watermarking
@@
As illustrated in Section II-C, MOVE [24] relies on a latent assumption that the model can sufficiently learn the embedded external dataset-specific features via modifying a few training samples with style transfer. However, since MOVE is designed for victim models trained from scratch, this assumption may not hold when fine-tuning is applied. Therefore, this method may fail to provide effective ownership verification. In this section, we demonstrate this limitation.
@@
Settings. We evaluate MOVE’s effectiveness on CIFAR-10 [54] using ResNet18 [55] under three settings: fine-tuning, unlearning-based fine-tuning, and training from scratch. Specifically, let $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}$ denote the benign training set, $\bm{x}_{s}$ be a defender-specified style image, and $T:\mathcal{X}\times\mathcal{X}\rightarrow\mathcal{X}$ be a pre-trained style transformer. Following the setting of MOVE [24], we randomly select $10\%$ samples ($i.e.$, $\mathcal{D}_{s}$) from $\mathcal{D}$ to generate their transformed versions $\mathcal{D}_{t}=\{(\bm{x}^{\prime},y)|\bm{x}^{\prime}=T(\bm{x},\bm{x}_{s}),(\bm{x},y)\in\mathcal{D}_{s}\}$. Under the original assumption of MOVE, the victim model should sufficiently learn the external features from $\mathcal{D}_{t}$ by training or fine-tuning on $\mathcal{D}_{b}\cup\mathcal{D}_{t}$. To evaluate this assumption under fine-tuning process, three ResNet18 models are then prepared: (1) ResNet18-FT: Fine-tuned from a pre-trained ResNet18 on $\mathcal{D}_{b}\cup\mathcal{D}_{t}$, where the pre-trained ResNet18 is trained with ImageNet [56] dataset and $\mathcal{D}_{b}\triangleq\mathcal{D}\backslash\mathcal{D}_{t}$. (2) ResNet18-UL: First unlearned on randomly relabeled $\mathcal{D}$ for 10 epochs before fine-tuning on $\mathcal{D}_{b}\cup\mathcal{D}_{t}$. (3) ResNet18-S: Trained from scratch on $\mathcal{D}_{b}\cup\mathcal{D}_{t}$.
@@
We evaluate the effectiveness of MOVE on ResNet18-FT, ResNet18-UL, and ResNet18-S under the simplest form of model stealing, where the victim model is directly copied and fine-tuned to create the stolen models. Following the setting in MOVE [24], we adopt the p-value as the evaluation metric, where the smaller the p-value, the more confident that MOVE believes the model stealing happened. We mark failed verification cases ($i.e.$, p-value $>0.01$ for stolen models) in red. We also test the accuracy of each model on the transformed dataset $\mathcal{D}_{t}$ and the unmodified CIFAR-10 test dataset to assess whether they have learned the external features in $\mathcal{D}_{t}$.
TABLE I: Failure of MOVE towards the fine-tuned model.
@@
Fine-tuning 0.48 $10^{-3}$ $10^{-10}$
TABLE II: Accuracy on $\mathcal{D}_{t}$ and the benign test dataset.
ResNet18-FT ResNet18-UL ResNet18-S
@@
Test Dataset 93.16% 87.50% 92.07%
Results. As shown in Table I, MOVE successfully detects model stealing for ResNet18-UL and ResNet18-S, while it fails to verify ownership for the fine-tuned model ResNet18-FT. This performance gap correlates with their learning capabilities of the external features. As shown in Table II, ResNet18-FT achieves only 17.54% accuracy on the transformed dataset $\mathcal{D}_{t}$, compared to 42.45% for ResNet18-UL and 88.17% for ResNet18-S. These findings reveal two key insights: First, fine-tuning process preserves insufficient capacity for learning external dataset-specific features, likely due to interference from the pre-trained model’s existing knowledge. Second, the unlearning process in ResNet18-UL appears to mitigate this interference, enabling better acquisition of external features. The results collectively demonstrate that MOVE’s verification mechanism, which relies on learning external features as dataset-specific features, is ineffective for fine-tuning scenarios due to interference from learned knowledge in the pre-trained foundation model.
III-C Revisiting Model Fingerprinting
As discussed in Section II-C, model fingerprinting methods are primarily designed for simple model stealing attacks and show limited robustness against advanced stealing methods [25, 24], such as knowledge distillation [11, 13]. Experiment results demonstrating these limitations are presented in Section V. In this section, we focus specifically on the false positive problem in model fingerprinting, where benign models may be incorrectly identified as stolen.
Specifically, we take dataset inference (DI) [25], a representative fingerprinting method, as the study case. As demonstrated in previous studies [50, 24], when verifying the ownership of models trained from scratch, DI easily leads to misjudgment when the training set of the suspicious models have a similar distribution to the victim model training set. As illustrated in the following experiments, this limitation also applies to the fine-tuned victim model.
Settings. We conduct experiments on the CIFAR-10 [54] dataset with VGG16 [57] and ResNet18 [55]. Specifically, we randomly split the original training set $\mathcal{D}$ into two disjoint subsets, $\mathcal{D}_{l}$ and $\mathcal{D}_{r}$. We use the pre-trained VGG16 and ResNet18 models, both trained on the ImageNet [56] dataset, as the foundation models. Three different models are then prepared: (1) VGG16-FT-$\mathcal{D}_{l}$: Fine-tuned from the pre-trained VGG16 on $\mathcal{D}_{l}$. (2) ResNet18-FT-$\mathcal{D}_{r}$: Fine-tuned from the pre-trained ResNet18 on $\mathcal{D}_{r}$. (3) VGG16-FT-$\mathcal{D}_{l}^{\prime}$: Fine-tuned from the pre-trained VGG16 on a noisy dataset $\mathcal{D}_{l}^{\prime}\triangleq\{(\bm{x}^{\prime},y)|\bm{x}^{\prime}=\bm{x}+\mathcal{N}(0,16),(\bm{x},y)\in\mathcal{D}_{l}\}$.
In the model ownership verification process, we verify whether the VGG16-FT-$\mathcal{D}_{l}$ and VGG16-FT-$\mathcal{D}_{l}^{\prime}$ are stolen from ResNet18-FT-$\mathcal{D}_{r}$, and whether ResNet18-FT-$\mathcal{D}_{r}$ is stolen from VGG16-FT-$\mathcal{D}_{l}$. Following the setting of dataset inference [25], we adopt the p-value as the evaluation metric, where the smaller the p-value, the more confident that dataset inference believes the model stealing happened.
Results. As shown in Table III, dataset inference misjudges in all cases. Specifically, ResNet18-FT-$\mathcal{D}_{r}$, VGG16-FT-$\mathcal{D}_{l}$, and VGG16-FT-$\mathcal{D}_{l}^{\prime}$ are all fine-tuned on different datasets, which means they should not be considered as stolen models of one another. However, dataset inference misjudges ResNet18-FT-$\mathcal{D}_{r}$ as stolen from VGG16-FT-$\mathcal{D}_{l}$ and misclassifies VGG16-FT-$\mathcal{D}_{l}$ and VGG16-FT-$\mathcal{D}_{l}^{\prime}$ as stolen from ResNet18-FT-$\mathcal{D}_{r}$. These misjudgments are probably because the fine-tuned models learn similar common features from $\mathcal{D}_{r}$, $\mathcal{D}_{l}$, and $\mathcal{D}_{l}^{\prime}$.
TABLE III: Misjudgments of DI towards fine-tuned models.
ResNet18-FT-$\mathcal{D}_{r}$ VGG16-FT-$\mathcal{D}_{l}$ VGG16-FT-$\mathcal{D}_{l}^{\prime}$
Accuary 92.45% 92.94% 63.53%
@@ III-C Revisiting Model Fingerprinting
IV The Proposed Method
Figure 2: The main pipeline of Holmes. Step 1. Creating Shadow Models: This step involves generating two shadow models to represent dataset-specific features. (a) The poisoned shadow model disrupts dataset-specific features while preserving common features through backdoor attack. (b) The benign shadow model introduces distinct dataset-specific features by fine-tuning on a filtered dataset with similar common features. Step 2. Training Ownership Meta-Classifier: A meta-classifier is trained using the output differences between the shadow models and the victim model to verify ownership. Step 3. Ownership Verification with Hypothesis Test: The final step involves conducting model ownership verification through a hypothesis test to mitigate randomness and further enhance robustness. The verification process requires no modifications to the victim model, ensuring the harmlessness of the proposed method.
@@ IV The Proposed Method
Based on the understandings in Section III, we propose Holmes, which exploits inherent yet dataset-specific features for verification via decoupling the common features. As shown in Figure 2, our method consists of three main stages, including (1) creating shadow models, (2) training ownership meta-classifier, and (3) ownership verification with hypothesis test. The pseudocode of Holmes is in the appendix. The technical details are in the following subsections.
IV-A Creating Shadow Models
@@
The Poisoned Shadow Model. As illustrated in Section III, similar common features may lead to misjudgment. If we can erase the impact of common features, the misjudgment may be avoided. However, defining these features is challenging due to the complex learning dynamics of LVMs. Luckily, we can decouple common features in an indirect way.
@@
We hereby propose to create a poisoned shadow model $P$ that retains common features learned by the victim model, while its dataset-specific features are disrupted. Thus, we can indirectly capture the dataset-specific features while decoupling the common features via the output difference between $V$ and $P$. The construction of $P$ is achieved via poisoning the most well-learned personalization samples (with the lowest loss value) with the label-inconsistent backdoor attack ($e.g.$, BadNets [28]). Specifically, for the training set $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}$, we select $\gamma\%$ samples with the lowest loss values (dubbed $\mathcal{D}_{s}$), where the loss function is cross-entropy. We choose these samples since they correspond to the model’s propensity to memorize dataset-specific patterns (instead of common features) [29]. We utilize the label-inconsistent backdoor attack to poison $\mathcal{D}_{s}$, since backdoor attack allows a high performance of poisoned models in predicting benign testing samples, thus preserving common features. Specifically, the defender first defines the target class $y_{t}$, trigger pattern $\bm{t}$, and a pre-defined poisoned image generator $G(\cdot)$ for backdoor attack. The poisoned subset $\mathcal{D}_{p}$ is created as $\mathcal{D}_{p}=\{(\bm{x}^{\prime},y_{t})|\bm{x}^{\prime}=G(\bm{x};\bm{t}),(\bm{x},y)\in\mathcal{D}_{s}\}$. The shadow model $P$ is created via fine-tuning $V$ with $\mathcal{D}_{p}\cup\mathcal{D}_{b}$, where $\mathcal{D}_{b}\triangleq\mathcal{D}\backslash\mathcal{D}_{s}$. Since $P$ maintains $V$’ common features while disrupting its dataset-specific features, we can capture the dataset-specific features by the output difference between $V$ and $P$ on $\mathcal{D}_{s}$ (to a large extent).
@@
The Benign Shadow Model. In order to further mitigate false positives from independent models with related dataset-specific features, we introduce the benign shadow $B$ model for reference. Generally, $B$ exhibits dataset-specific features that are similar but distinct from those of the victim model. Thus, the output difference between $B$ and $P$ reveals dataset-specific features that are distinct from those of $V$. To construct $B$, we fine-tune the same foundation model $F$ on a filtered dataset $\mathcal{D}_{f}$ which excludes $\lambda\%$ of the $V$’s training samples with the lowest loss values. By dropping these well-learned samples, $\mathcal{D}_{f}$ contains data that encourages $B$ to develop distinct dataset-specific features while preserving common features shared with $V$. Thus, the output difference between $B$ and $P$ highlights dataset-specific features distinct from those revealed by the output difference between $V$ and $P$.
IV-B Training Ownership Meta-Classifier

### v2

@@ 3.1 Preliminaries
where $\mathcal{L}(\cdot)$ is the loss function ($i.e.$, cross-entropy), and $\bm{\theta}$ represents the parameters of the foundation model $F$. This fine-tuning process adapts $F$ to the specific downstream task defined by $\mathcal{D}$, rather than trains the model from scratch.
@@
For a model ownership verification method to be effective in practice, it must satisfy two key requirements: effectiveness and harmlessness. Effectiveness means that the defense must accurately determine whether the suspicious model has been stolen from the victim, regardless of the model stealing technique employed. At the same time, it should not misclassify independently trained or fine-tuned benign models as stolen models. Harmlessness ensures that the defense does not introduce additional security risks, such as backdoors. Furthermore, if the method modifies the training set, harmlessness implies that models trained or fine-tuned with the modified dataset should exhibit behavior similar to those trained with the original dataset. In particular, this paper adopts a stricter harmlessness criterion, requiring that the defense must not modify the victim model’s parameters in any form, including direct alteration, additional training, or fine-tuning.
@@
Threat Model. In this paper, we consider the black-box setting for model ownership verification. Specifically, we assume that the defenders can only query and obtain the predicted probabilities (logit outputs) from the suspicious model, with no access to model source files or parameters, intermediate computational results ($e.g.$, gradients), or any information about the model stealing process.
3.2 Revisiting Harmless Model Watermarking
@@
As illustrated in Section 2.3, MOVE [li2025move] relies on a latent assumption that the model can sufficiently learn the embedded external dataset-specific features via modifying a few training samples with style transfer. However, since MOVE is designed for victim models trained from scratch, this assumption may not hold when fine-tuning is applied. Therefore, this method may fail to provide effective ownership verification. In this section, we demonstrate this limitation.
@@
Settings. We evaluate MOVE’s effectiveness on CIFAR-10 [krizhevsky2009learning] using ResNet18 [he2016deep] under three settings: fine-tuning, unlearning-based fine-tuning, and training from scratch. Specifically, let $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}$ denote the benign training set, $\bm{x}_{s}$ be a defender-specified style image, and $T:\mathcal{X}\times\mathcal{X}\rightarrow\mathcal{X}$ be a pre-trained style transformer. Following the setting of MOVE [li2025move], we randomly select $10\%$ samples ($i.e.$, $\mathcal{D}_{s}$) from $\mathcal{D}$ to generate their transformed versions $\mathcal{D}_{t}=\{(\bm{x}^{\prime},y)|\bm{x}^{\prime}=T(\bm{x},\bm{x}_{s}),(\bm{x},y)\in\mathcal{D}_{s}\}$. Under the original assumption of MOVE, the victim model should sufficiently learn the external features from $\mathcal{D}_{t}$ by training or fine-tuning on $\mathcal{D}_{b}\cup\mathcal{D}_{t}$. To evaluate this assumption under fine-tuning process, three ResNet18 models are then prepared: (1) ResNet18-FT: Fine-tuned from a pre-trained ResNet18 on $\mathcal{D}_{b}\cup\mathcal{D}_{t}$, where the pre-trained ResNet18 is trained with ImageNet [deng2009imagenet] dataset and $\mathcal{D}_{b}\triangleq\mathcal{D}\backslash\mathcal{D}_{t}$. (2) ResNet18-UL: First unlearned on randomly relabeled $\mathcal{D}$ for 10 epochs before fine-tuning on $\mathcal{D}_{b}\cup\mathcal{D}_{t}$. (3) ResNet18-S: Trained from scratch on $\mathcal{D}_{b}\cup\mathcal{D}_{t}$.
@@
We evaluate the effectiveness of MOVE on ResNet18-FT, ResNet18-UL, and ResNet18-S under the simplest form of model stealing, where the victim model is directly copied and fine-tuned to create the stolen models. Following the setting in MOVE [li2025move], we adopt the p-value as the evaluation metric, where the smaller the p-value, the more confident that MOVE believes the model stealing happened. We mark failed verification cases ($i.e.$, p-value >0.01>0.01 for stolen models) in red. We also test the accuracy of each model on the transformed dataset $\mathcal{D}_{t}$ and the unmodified CIFAR-10 test dataset to assess whether they have learned the external features in $\mathcal{D}_{t}$.
Table 1: Failure of MOVE on fine-tuned models.
@@
Fine-tuning 0.48 $^{-3} 1010$$^{-10}$
Table 2: Accuracy on $\mathcal{D}_{t}$ and the benign set.
Test Set $\downarrow$ ResNet18-FT ResNet18-UL ResNet18-S
@@
Benign Set 93.16% 87.50% 92.07%
Results. As shown in Table 1, MOVE successfully detects model stealing for ResNet18-UL and ResNet18-S, while it fails to verify ownership for the fine-tuned model ResNet18-FT. This performance gap correlates with their learning capabilities of the external features. As shown in Table 2, ResNet18-FT achieves only 17.54% accuracy on the transformed dataset $\mathcal{D}_{t}$, compared to 42.45% for ResNet18-UL and 88.17% for ResNet18-S. These findings reveal two key insights: First, fine-tuning process lacks sufficient capacity for learning external dataset-specific features, likely due to interference from the pre-trained model’s existing knowledge. Second, the unlearning process in ResNet18-UL appears to mitigate this interference, enabling better acquisition of external features. The results collectively demonstrate that MOVE’s verification mechanism, which relies on learning external features as dataset-specific features, is ineffective for fine-tuning scenarios due to interference from learned knowledge in the pre-trained foundation model.
Table 3: Misjudgments of DI on fine-tuned models.
ResNet18-FT-$\mathcal{D}_{r}$ $\mathcal{D}_{l} VGG16-FT$-$\mathcal{D}_{l}^{\prime}$
Accuracy 92.45% 92.94% 63.53%
@@
3.3 Revisiting Model Fingerprinting
As discussed in Section 2.3, model fingerprinting methods are primarily designed for simple model stealing attacks and show limited robustness against advanced stealing methods [maini2021dataset, li2025move], such as knowledge distillation [li2023curriculum, sun2024logit]. Experiment results demonstrating these limitations are presented in Section 5. In this section, we focus specifically on the false positive problem in model fingerprinting, where benign models may be incorrectly identified as stolen.
Specifically, we take dataset inference (DI) [maini2021dataset], a representative fingerprinting method, as the study case. As demonstrated in previous studies [li2022defending, li2025move], when verifying the ownership of models trained from scratch, DI easily leads to misjudgment when the training set of the suspicious models has a similar distribution to the victim model training set. As illustrated in the following experiments, this limitation also applies to the fine-tuned victim models.
Settings. We conduct experiments on the CIFAR-10 [krizhevsky2009learning] dataset with VGG16 [simonyan2014very] and ResNet18 [he2016deep]. Specifically, we randomly split the original training set $\mathcal{D}$ into two disjoint subsets, $\mathcal{D}_{l}$ and $\mathcal{D}_{r}$. We use the pre-trained VGG16 and ResNet18 models, both trained on the ImageNet [deng2009imagenet], as the foundation models. Three different models are then prepared: (1) VGG16-FT-$\mathcal{D}_{l}$: Fine-tuned from the pre-trained VGG16 on $\mathcal{D}_{l}$. (2) ResNet18-FT-$\mathcal{D}_{r}$: Fine-tuned from the pre-trained ResNet18 on $\mathcal{D}_{r}$. (3) VGG16-FT-$\mathcal{D}_{l}^{\prime}$: Fine-tuned from the pre-trained VGG16 on a noisy dataset $\mathcal{D}_{l}^{\prime}\triangleq\{(\bm{x}^{\prime},y)|\bm{x}^{\prime}=\bm{x}+\mathcal{N}(0,16),(\bm{x},y)\in\mathcal{D}_{l}\}$.
In the model ownership verification process, we verify whether VGG16-FT-$\mathcal{D}_{l}$ and VGG16-FT-$\mathcal{D}_{l}^{\prime}$ are stolen from ResNet18-FT-$\mathcal{D}_{r}$, and whether ResNet18-FT-$\mathcal{D}_{r}$ is stolen from VGG16-FT-$\mathcal{D}_{l}$. Following the setting of dataset inference [maini2021dataset], we adopt the p-value as the evaluation metric, where the smaller the p-value, the more confident that dataset inference believes the model stealing happened. It is worth noting that $\mathcal{D}_{l}$, $\mathcal{D}_{r}$, and $\mathcal{D}_{l}^{\prime}$ are highly similar in distribution but not identical, and the three models are not involved in any model stealing between each other.
Results. As shown in Table 3, DI misjudges all cases. Although ResNet18-FT-$\mathcal{D}_{r}$, VGG16-FT-$\mathcal{D}_{l}$, and VGG16-FT-$\mathcal{D}_{l}^{\prime}$ are fine-tuned on different datasets and should not be considered as stolen, the method incorrectly claims ownership between them. These misjudgments are probably because the fine-tuned models learn similar common features from $\mathcal{D}_{r}$, $\mathcal{D}_{l}$, and $\mathcal{D}_{l}^{\prime}$.
4 The Proposed Method
@@
Based on the understandings in Section 3, where we revealed that existing harmless MOV methods either fail for fine-tuned models due to interference from pre-trained knowledge or suffer from misjudgments caused by shared common features, we propose Holmes, which exploits inherent yet dataset-specific features for verification via decoupling the common features. As shown in Figure 2, our method consists of three main stages, including (1) creating shadow models, (2) training ownership meta-classifier, and (3) ownership verification with hypothesis test. The pseudocode of Holmes is in Appendix B. The technical details are in the following subsections.
Figure 2: The main pipeline of Holmes. Step 1. Creating Shadow Models: This step involves generating two shadow models to represent dataset-specific features. (a) The poisoned shadow model disrupts dataset-specific features while preserving common features through backdoor attack. (b) The benign shadow model introduces distinct dataset-specific features by fine-tuning on a filtered dataset with similar common features. Step 2. Training Ownership Meta-Classifier: A meta-classifier is trained using the output differences between the shadow models and the victim model to verify ownership. Step 3. Ownership Verification with Hypothesis Test: The final step involves conducting model ownership verification through a hypothesis test to mitigate randomness and further enhance robustness. The verification process requires no modifications to the victim model, ensuring the harmlessness of Holmes.
4.1 Creating Shadow Models
@@
The Poisoned Shadow Model. As illustrated in Section 3, similar common features may lead to misjudgment. If we can erase the impact of common features, the misjudgment may be avoided. However, defining these features is challenging due to the complex learning dynamics of LVMs. Luckily, we can decouple common features indirectly.
@@
We hereby propose to create a poisoned shadow model $P$ that retains common features learned by the victim model, while its dataset-specific features are disrupted. Thus, we can indirectly capture the dataset-specific features while decoupling the common features via the output difference between $V$ and $P$. $P$ is constructed by poisoning the most well-learned personalization samples (with the lowest loss values) with the label-inconsistent backdoor attack ($e.g.$, BadNets [gu2019badnets]). We choose these samples since they correspond to the model’s propensity to memorize dataset-specific patterns (instead of common features) [ilyas2019adversarial]. Specifically, for the training set $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}$, we select $\gamma\%$ samples with the lowest loss values (dubbed $\mathcal{D}_{s}$), where the loss function is cross-entropy. We adopt the label-inconsistent backdoor attack to poison $\mathcal{D}_{s}$, as this type of attack allows a high performance of poisoned models in predicting benign samples, thus preserving common features. Specifically, the defender first defines the target class $y_{t}$, trigger pattern $\bm{t}$, and a poisoned image generator $G(\cdot)$. The poisoned subset $\mathcal{D}_{p}$ is created as $\mathcal{D}_{p}=\{(\bm{x}^{\prime},y_{t})|\bm{x}^{\prime}=G(\bm{x};\bm{t}),(\bm{x},y)\in\mathcal{D}_{s}\}$. $P$ is created via fine-tuning $V$ with $\mathcal{D}_{p}\cup\mathcal{D}_{b}$, where $\mathcal{D}_{b}\triangleq\mathcal{D}\backslash\mathcal{D}_{s}$. Since $P$ maintains $V$’s common features while disrupting its dataset-specific features, we can capture the dataset-specific features by the output difference between $V$ and $P$ on $\mathcal{D}_{s}$ indirectly.
@@
The Benign Shadow Model. In order to further mitigate false positives from independent models with related dataset-specific features, we introduce the benign shadow model $B$ for reference. Generally, $B$ exhibits dataset-specific features that are similar but distinct from those of the victim model. Thus, the output difference between $B$ and $P$ reveals dataset-specific features that are distinct from those of $V$. To construct $B$, we fine-tune the same foundation model $F$ on a filtered dataset $\mathcal{D}_{f}$ which excludes $\lambda\%$ of the $V$’s training samples with the lowest loss values. By dropping these well-learned samples, $\mathcal{D}_{f}$ contains data that encourages $B$ to develop distinct dataset-specific features while preserving common features shared with $V$. Thus, the output difference between $B$ and $P$ highlights dataset-specific features distinct from those revealed by the output difference between $V$ and $P$.
4.2 Training Meta-Classifier